# Accelerating Ill-conditioned Hankel Matrix RecoveryHanQin Cai, Longxiu Huang, Xiliang Lu, and Juntao You

### v1

@@ 5.1 Technical lemmas
(18) $\left\lVert{\bm{L}}_{k}{\bm{R}}_{k}^{*}-\mathcal{G}\bm{z}_{\star}\right\lVert_{\mathrm{F}}=$ $~{}\left\lVert\bm{\Delta}_{{\bm{L}}_{k}}\bm{R}_{\star}^{*}+\bm{L}_{\star}\bm{\Delta}_{{\bm{R}}_{k}}^{*}+\bm{\Delta}_{{\bm{L}}_{k}}\bm{\Delta}_{{\bm{R}}_{k}}^{*}\right\lVert_{\mathrm{F}}$
(19) $\leq$ $~{}\left\lVert\bm{\Delta}_{{\bm{L}}_{k}}\bm{\Sigma}_{\star}^{1/2}{\bm{V}}_{\star}^{*}\right\lVert_{\mathrm{F}}+\left\lVert{\bm{U}}_{\star}\bm{\Sigma}_{\star}^{1/2}\bm{\Delta}_{{\bm{R}}_{k}}^{*}\right\lVert_{\mathrm{F}}+\left\lVert\bm{\Delta}_{{\bm{L}}_{k}}\bm{\Delta}_{{\bm{R}}_{k}}^{*}\right\lVert_{\mathrm{F}}$
@@
(20) $\leq$ $~{}(1+\frac{\varepsilon_{0}}{2})\left(\left\lVert\bm{\Delta}_{{\bm{L}}_{k}}\bm{\Sigma}_{\star}^{1/2}\right\lVert_{\mathrm{F}}+\left\lVert\bm{\Delta}_{{\bm{R}}_{k}}\bm{\Sigma}_{\star}^{1/2}\right\lVert_{\mathrm{F}}\right).$
@@
Thus, by noticing $\left(\left\lVert\bm{\Delta}_{{\bm{L}}_{k}}\bm{\Sigma}_{\star}^{1/2}\right\lVert_{\mathrm{F}}+\left\lVert\bm{\Delta}_{{\bm{R}}_{k}}\bm{\Sigma}_{\star}^{1/2}\right\lVert_{\mathrm{F}}\right)^{2}\leq 2d_{k}^{2}$, we complete the proof.
5.3 Proof of Theorem 3.3 (guaranteed initialization)
Proof 5.15.
@@
This proof is under event (11) and we bound $d_{0}$ in the following. Following the notation in Algorithm 1, we denote
(21) $\bm{M}_{0}:=p^{-1}\mathcal{G}\left(\Pi_{\Omega}\bm{f}-\bm{s}_{0}\right),\quad\text{and}~{}\bm{M}_{0}^{(r)}~{}\text{be the top-$r$ SVD of}~{}\bm{M}_{0}.$
@@
For some universal constant $\tilde{c}_{0}$, if provided $m\geq 64\tilde{c}_{0}^{2}\varepsilon_{0}^{-2}\kappa^{2}c_{s}\mu r^{2}\log n$, we have
(22) $\left\lVert\bm{M}_{0}-\mathcal{G}\bm{z}_{\star}\right\lVert_{2}$ $\leq\left\lVert\bm{M}_{0}-p^{-1}\mathcal{G}\Pi_{\Omega}\bm{z}_{\star}\right\lVert_{2}+\left\lVert p^{-1}\mathcal{G}\Pi_{\Omega}\bm{z}_{\star}-\mathcal{G}\bm{z}_{\star}\right\lVert_{2}$
(23) $=\left\lVert p^{-1}\mathcal{G}\left(\Pi_{\Omega}\bm{s}_{\star}-\bm{s}_{0}\right)\right\lVert_{2}+\left\lVert\left(p^{-1}\mathcal{G}\Pi_{\Omega}-\mathcal{G}\right)\bm{z}_{\star}\right\lVert_{2}$
@@
(24) $\leq 2\alpha n\big{\|}\mathcal{G}\big{(}\Pi_{\Omega}\bm{s}_{\star}-\bm{s}_{0}\big{)}\big{\|}_{\infty}+\tilde{c}_{0}\sqrt{(\mu c_{s}r\log n)/m}\sigma^{\star}_{1}$
(25) $\leq 4\alpha n\big{\|}\mathcal{G}\bm{z}_{\star}\big{\|}_{\infty}+0.125\varepsilon_{0}\sigma^{\star}_{r}r^{-\frac{1}{2}}$
(26) $\leq 4\alpha c_{s}\mu r\sigma^{\star}_{1}+0.125\varepsilon_{0}\sigma^{\star}_{r}r^{-\frac{1}{2}},$
@@
where the second step follows from $\|\mathcal{G}\left(\Pi_{\Omega}\bm{f}-\bm{s}_{0}\right)-\mathcal{G}\Pi_{\Omega}\bm{z}_{\star}\|_{2}=\|\mathcal{G}\left(\Pi_{\Omega}\bm{s}_{\star}-\bm{s}_{0}\right)\|_{2}$. The third line follows from [2, Lemma 6] (as $\Pi_{\Omega}\bm{s}_{\star}-\bm{s}_{0}$ is at most $2\alpha m$-sparse) and Lemma 5.1. The fourth line follows from (14) with $\bm{z}=\bm{0}$, i.e., $\left\lVert\mathcal{G}\left(\Pi_{\Omega}\bm{s}_{\star}-\bm{s}_{0}\right)\right\lVert_{\infty}=\left\lVert\mathcal{W}\left(\Pi_{\Omega}\bm{s}_{\star}-\bm{s}_{0}\right)\right\lVert_{\infty}\leq 2\left\lVert\mathcal{W}\Pi_{\Omega}\bm{z}_{\star}\right\lVert_{\infty}=2\left\lVert\mathcal{G}\Pi_{\Omega}\bm{z}_{\star}\right\lVert_{\infty}$. And the last inequality follows from
@@
(27) $\left\lVert\bm{M}_{0}^{(r)}-\mathcal{G}\bm{z}_{\star}\right\lVert_{2}$ $\leq\left\lVert\bm{M}_{0}^{(r)}-\bm{M}_{0}\right\lVert_{2}+\left\lVert\bm{M}_{0}-\mathcal{G}\bm{z}_{\star}\right\lVert_{2}$
(28) $\leq 2\left\lVert\bm{M}_{0}-\mathcal{G}\bm{z}_{\star}\right\lVert_{2}~{}\leq 8\alpha c_{s}\mu r\kappa\sigma^{\star}_{r}+0.25\varepsilon_{0}\sigma^{\star}_{r}r^{-\frac{1}{2}},$
where the second inequality follows from the definition of $\bm{M}_{0}^{(r)}$ and Eckart-Young-Mirsky theorem. Then, by [33, Lemma 24] we have
@@
Then, for any given $\varepsilon_{0}\in(0,1)$, we have $\mathrm{dist}\left({\bm{U}}_{0}\bm{\Sigma}_{0}^{1/2},{\bm{V}}_{0}\bm{\Sigma}_{0}^{1/2};\bm{L}_{\star},\bm{R}_{\star}\right)\leq\varepsilon_{0}\sigma^{\star}_{r}<\sigma^{\star}_{r}$ provided $0\leq\alpha\leq\frac{\varepsilon_{0}}{50c_{s}\kappa\mu r\sqrt{r}}$. Thus, for $c\geq 1+\varepsilon_{0}$, by Lemma 5.12 we have
@@
and $\max\{\left\lVert{\bm{L}}_{0}{\bm{R}}_{0}^{*}\right\lVert_{2,\infty},\left\lVert{\bm{R}}_{0}{\bm{L}}_{0}^{*}\right\lVert_{2,\infty}\}\leq c\sqrt{\frac{c_{s}\mu r}{n}}\sigma^{\star}_{1}$. This completes the proof.
5.4 Proof of Theorems 3.5 and 3.8 (local convergence)
The proof follows a similar route established in [33], but the techniques and details are quite involved due to the extra Hankel structure, as well as the simultaneous presence of outliers and missing data.
@@
(29a) $\max\{\left\lVert\bm{\Delta}_{{\bm{L}}_{k}}\bm{\Sigma}_{\star}^{-1/2}\right\lVert_{2},\left\lVert\bm{\Delta}_{{\bm{R}}_{k}}\bm{\Sigma}_{\star}^{-1/2}\right\lVert_{2}\}$ $\leq\varepsilon_{0},$
(29b) $\max\left\{\left\lVert\bm{L}_{\natural}\bm{\Sigma}_{\star}^{1/2}\right\lVert_{2,\infty},\left\lVert\bm{R}_{\natural}\bm{\Sigma}_{\star}^{1/2}\right\lVert_{2,\infty}\right\}$ $\leq\frac{c}{1-\varepsilon_{0}}\sqrt{\frac{c_{s}\mu r}{n}}\sigma^{\star}_{1},$
(29c) $\max\left\{\left\lVert\bm{L}_{\natural}\bm{\Sigma}_{\star}^{-1/2}\right\lVert_{2,\infty},\left\lVert\bm{R}_{\natural}\bm{\Sigma}_{\star}^{-1/2}\right\lVert_{2,\infty}\right\}$ $\leq\frac{c\kappa}{1-\varepsilon_{0}}\sqrt{\frac{c_{s}\mu r}{n}},$
(29d) $\left\lVert\bm{R}_{\natural}\left(\bm{R}_{\natural}^{*}\bm{R}_{\natural}\right)^{-1}\bm{\Sigma}_{\star}^{1/2}\right\lVert_{2}$ $\leq\frac{1}{1-\varepsilon_{0}}.$
Proof 5.17.
@@
then by $\left\lVert{\bm{L}}_{k}{\bm{R}}_{k}^{*}\right\lVert_{2,\infty}\geq\sigma_{r}(\bm{R}_{\natural}\bm{\Sigma}_{\star}^{-1/2})\left\lVert\bm{L}_{\natural}\bm{\Sigma}_{\star}^{1/2}\right\lVert_{2,\infty}$ we have $\left\lVert\bm{L}_{\natural}\bm{\Sigma}_{\star}^{1/2}\right\lVert_{2,\infty}\leq\frac{\left\lVert{\bm{L}}_{k}{\bm{R}}_{k}^{*}\right\lVert_{2,\infty}}{1-\varepsilon_{0}}$. Similarly, we have $\left\lVert\bm{R}_{\natural}\bm{\Sigma}_{\star}^{1/2}\right\lVert_{2,\infty}\leq\frac{\left\lVert{\bm{L}}_{k}{\bm{R}}_{k}^{*}\right\lVert_{2,\infty}}{1-\varepsilon_{0}}$ and conclude (29b). By using $\left\lVert{\bm{A}}{\bm{B}}\right\lVert_{2,\infty}\leq\left\lVert{\bm{A}}\right\lVert_{2,\infty}\left\lVert{\bm{B}}\right\lVert_{2}$, (29c) follows directly from (29b). (29d) follows from [33, Lemma 25] and (29a).
Lemma 5.18.
For all matrices $({\bm{L}}_{k},{\bm{R}}_{k})\in\mathcal{E}(\varepsilon_{0}\sigma^{\star}_{r},c\sqrt{\frac{c_{s}\mu r}{n}}\sigma^{\star}_{1})$, it holds
@@
$\leq 2\sqrt{p}(1+\varepsilon_{0})\left(\left\lVert\bm{\Delta}_{{\bm{L}}_{k}}\bm{\Sigma}_{\star}^{1/2}\right\lVert_{\mathrm{F}}+\left\lVert\bm{\Delta}_{{\bm{R}}_{k}}\bm{\Sigma}_{\star}^{1/2}\right\lVert_{\mathrm{F}}\right)$
@@
where the first step follows from $\left\lvert\left\langle\Pi_{\Omega}(\bm{u}+\bm{v}),\bm{u}+\bm{v}\right\rangle\right\rvert^{\frac{1}{2}}\leq\left\lvert\left\langle\Pi_{\Omega}\bm{u},\bm{u}\right\rangle\right\rvert^{\frac{1}{2}}+\left\lvert\left\langle\Pi_{\Omega}\bm{v},\bm{v}\right\rangle\right\rvert^{\frac{1}{2}}$ for $\bm{u},\bm{v}\in\mathbb{C}^{m}$, the second inequality follows from Lemma 5.8, and in the last inequality we have used $\left\lVert\bm{\Delta}_{{\bm{L}}_{k}}\bm{\Sigma}_{\star}^{1/2}{\bm{V}}_{\star}^{*}\right\lVert_{\mathrm{F}}=\left\lVert\bm{\Delta}_{{\bm{L}}_{k}}\bm{\Sigma}_{\star}^{1/2}\right\lVert_{\mathrm{F}}$. For the last term, provided $m\geq 6\varepsilon_{0}^{-2}C_{\varepsilon_{0},\kappa}^{2}c_{s}^{2}\mu^{2}r^{2}\log n$, we have
$~{}p^{-1}\left\lvert\left\langle\Pi_{\Omega}\mathcal{G}^{*}\left(\bm{\Delta}_{{\bm{L}}_{k}}\bm{\Delta}_{{\bm{R}}_{k}}^{*}\right),\mathcal{G}^{*}\left(\bm{\Delta}_{{\bm{L}}_{k}}\bm{\Delta}_{{\bm{R}}_{k}}^{*}\right)\right\rangle\right\rvert$
@@
$~{}+\sqrt{24p^{-1}n\log n}\left\lVert\bm{\Delta}_{{\bm{L}}_{k}}\bm{\Sigma}_{\star}^{1/2}\right\lVert_{2,\infty}\left\lVert\bm{\Delta}_{{\bm{R}}_{k}}\bm{\Sigma}_{\star}^{-1/2}\right\lVert_{2,\infty}\left\lVert\bm{\Delta}_{{\bm{L}}_{k}}\bm{\Sigma}_{\star}^{1/2}\right\lVert_{\mathrm{F}}\left\lVert\bm{\Delta}_{{\bm{R}}_{k}}\bm{\Sigma}_{\star}^{-1/2}\right\lVert_{\mathrm{F}}$
@@
where the first inequality follows from Lemma 5.6, the second inequality follows from the inequalities $\left\lVert\bm{\Delta}_{{\bm{R}}_{k}}\bm{\Sigma}_{\star}^{-1/2}\right\lVert_{\mathrm{F}}=\left\lVert\bm{\Delta}_{{\bm{R}}_{k}}\bm{\Sigma}_{\star}^{1/2}\bm{\Sigma}_{\star}^{-1}\right\lVert_{\mathrm{F}}\leq\frac{1}{\sigma^{\star}_{r}}\left\lVert\bm{\Delta}_{{\bm{R}}_{k}}\bm{\Sigma}_{\star}^{1/2}\right\lVert_{\mathrm{F}}$, and (by (29c))
(30) $\begin{split}\left\lVert\bm{\Delta}_{{\bm{L}}_{k}}\bm{\Sigma}_{\star}^{-\frac{1}{2}}\right\lVert_{2,\infty}&\leq\left\lVert\bm{L}_{\natural}\bm{\Sigma}_{\star}^{-\frac{1}{2}}\right\lVert_{2,\infty}+\left\lVert\bm{L}_{\star}\bm{\Sigma}_{\star}^{-\frac{1}{2}}\right\lVert_{2,\infty}\leq\left(1+\frac{c\kappa}{1-\varepsilon_{0}}\right)\sqrt{\frac{c_{s}\mu r}{n}},\cr\left\lVert\bm{\Delta}_{{\bm{L}}_{k}}\bm{\Sigma}_{\star}^{\frac{1}{2}}\right\lVert_{2,\infty}&\leq\left\lVert\bm{L}_{\natural}\bm{\Sigma}_{\star}^{\frac{1}{2}}\right\lVert_{2,\infty}+\left\lVert\bm{L}_{\star}\bm{\Sigma}_{\star}^{\frac{1}{2}}\right\lVert_{2,\infty}\leq\left(1+\frac{c}{1-\varepsilon_{0}}\right)\sqrt{\frac{c_{s}\mu r}{n}}\sigma^{\star}_{1}.\end{split}$
@@
Putting all the bound of terms together, we conclude the lemma.
Lemma 5.20.

### v2

@@ 5.2 Proof of Proposition 1
$\left\lVert{\bm{L}}_{k}{\bm{R}}_{k}^{*}-\mathcal{G}\bm{z}_{\star}\right\lVert_{\mathrm{F}}=$ $~{}\left\lVert\bm{\Delta}_{{\bm{L}}_{k}}\bm{R}_{\star}^{*}+\bm{L}_{\star}\bm{\Delta}_{{\bm{R}}_{k}}^{*}+\bm{\Delta}_{{\bm{L}}_{k}}\bm{\Delta}_{{\bm{R}}_{k}}^{*}\right\lVert_{\mathrm{F}}$ (17)
$\leq$ $~{}\left\lVert\bm{\Delta}_{{\bm{L}}_{k}}\bm{\Sigma}_{\star}^{1/2}{\bm{V}}_{\star}^{*}\right\lVert_{\mathrm{F}}+\left\lVert{\bm{U}}_{\star}\bm{\Sigma}_{\star}^{1/2}\bm{\Delta}_{{\bm{R}}_{k}}^{*}\right\lVert_{\mathrm{F}}+\left\lVert\bm{\Delta}_{{\bm{L}}_{k}}\bm{\Delta}_{{\bm{R}}_{k}}^{*}\right\lVert_{\mathrm{F}}$ (18)
@@
$\leq$ $~{}(1+\frac{\varepsilon_{0}}{2})\left(\left\lVert\bm{\Delta}_{{\bm{L}}_{k}}\bm{\Sigma}_{\star}^{1/2}\right\lVert_{\mathrm{F}}+\left\lVert\bm{\Delta}_{{\bm{R}}_{k}}\bm{\Sigma}_{\star}^{1/2}\right\lVert_{\mathrm{F}}\right).$ (19)
@@
Thus, by noticing $\left(\left\lVert\bm{\Delta}_{{\bm{L}}_{k}}\bm{\Sigma}_{\star}^{1/2}\right\lVert_{\mathrm{F}}+\left\lVert\bm{\Delta}_{{\bm{R}}_{k}}\bm{\Sigma}_{\star}^{1/2}\right\lVert_{\mathrm{F}}\right)^{2}\leq 2d_{k}^{2}$, we complete the proof. ∎
5.3 Proof of Theorem 1 (guaranteed initialization)
Proof.
@@ 5.3 Proof of Theorem 1 (guaranteed initialization)
This proof is under event (10) and we bound $d_{0}$ in the following. Following the notation of Algorithm 1, we denote
$\bm{M}_{0}:=\frac{1}{p}\mathcal{G}\left(\Pi_{\Omega}\bm{f}-\bm{s}_{0}\right),\quad\text{and}~{}\bm{M}_{0}^{(r)}~{}\text{be the top-$r$ SVD of}~{}\bm{M}_{0}.$ (20)
@@
For some universal constant $\tilde{c}_{0}$, provided $m\geq 64\tilde{c}_{0}^{2}\varepsilon_{0}^{-2}\kappa^{2}c_{s}\mu r^{2}\log n$, we have
$\left\lVert\bm{M}_{0}-\mathcal{G}\bm{z}_{\star}\right\lVert_{2}$ $\leq\left\lVert\bm{M}_{0}-\frac{1}{p}\mathcal{G}\Pi_{\Omega}\bm{z}_{\star}\right\lVert_{2}+\left\lVert\frac{1}{p}\mathcal{G}\Pi_{\Omega}\bm{z}_{\star}-\mathcal{G}\bm{z}_{\star}\right\lVert_{2}$ (21)
$=\left\lVert\frac{1}{p}\mathcal{G}\left(\Pi_{\Omega}\bm{s}_{\star}-\bm{s}_{0}\right)\right\lVert_{2}+\left\lVert\left(\frac{1}{p}\mathcal{G}\Pi_{\Omega}-\mathcal{G}\right)\bm{z}_{\star}\right\lVert_{2}$ (22)
@@
$\leq 2\alpha n\big{\|}\mathcal{G}\big{(}\Pi_{\Omega}\bm{s}_{\star}-\bm{s}_{0}\big{)}\big{\|}_{\infty}+\tilde{c}_{0}\sqrt{(\mu c_{s}r\log n)/m}\sigma^{\star}_{1}$ (23)
$\leq 4\alpha n\big{\|}\mathcal{G}\bm{z}_{\star}\big{\|}_{\infty}+0.125\varepsilon_{0}\sigma^{\star}_{r}r^{-\frac{1}{2}}$ (24)
$\leq 4\alpha c_{s}\mu r\sigma^{\star}_{1}+0.125\varepsilon_{0}\sigma^{\star}_{r}r^{-\frac{1}{2}},$ (25)
@@
where the second step follows from $\|\mathcal{G}\left(\Pi_{\Omega}\bm{f}-\bm{s}_{0}\right)-\mathcal{G}\Pi_{\Omega}\bm{z}_{\star}\|_{2}=\|\mathcal{G}\left(\Pi_{\Omega}\bm{s}_{\star}-\bm{s}_{0}\right)\|_{2}$. The third line follows from [2, Lemma 6] ( as $\mathcal{G}=\mathcal{H}\mathcal{W}$ and $\mathcal{W}(\Pi_{\Omega}\bm{s}_{\star}-\bm{s}_{0})$ is at most $2\alpha m$-sparse) and Lemma 1. The fourth line follows from (13) with $\bm{z}=\bm{0}$, i.e., $\left\lVert\mathcal{G}\left(\Pi_{\Omega}\bm{s}_{\star}-\bm{s}_{0}\right)\right\lVert_{\infty}=\left\lVert\mathcal{W}\left(\Pi_{\Omega}\bm{s}_{\star}-\bm{s}_{0}\right)\right\lVert_{\infty}\leq 2\left\lVert\mathcal{W}\Pi_{\Omega}\bm{z}_{\star}\right\lVert_{\infty}=2\left\lVert\mathcal{G}\Pi_{\Omega}\bm{z}_{\star}\right\lVert_{\infty}$, because $\left\lVert\mathcal{G}\bm{z}\right\lVert_{\infty}$ and $\left\lVert\mathcal{W}\bm{z}\right\lVert_{\infty}$ with $\mathcal{W}\bm{z}=\bm{x}$ both equal the largest entry-wise magnitude in $\bm{x}$, see (1). And the last inequality follows from
@@
$\left\lVert\bm{M}_{0}^{(r)}-\mathcal{G}\bm{z}_{\star}\right\lVert_{2}$ $\leq\left\lVert\bm{M}_{0}^{(r)}-\bm{M}_{0}\right\lVert_{2}+\left\lVert\bm{M}_{0}-\mathcal{G}\bm{z}_{\star}\right\lVert_{2}$ (26)
$\leq 2\left\lVert\bm{M}_{0}-\mathcal{G}\bm{z}_{\star}\right\lVert_{2}~{}\leq 8\alpha c_{s}\mu r\kappa\sigma^{\star}_{r}+0.25\varepsilon_{0}\sigma^{\star}_{r}r^{-\frac{1}{2}},$ (27)
where the second inequality follows from the definition of $\bm{M}_{0}^{(r)}$ and Eckart-Young-Mirsky theorem. Then, by [36, Lemma 24] we have
@@
Then, for any given $\varepsilon_{0}\in(0,1)$, we have $\mathrm{dist}\left({\bm{U}}_{0}\bm{\Sigma}_{0}^{1/2},{\bm{V}}_{0}\bm{\Sigma}_{0}^{1/2};\bm{L}_{\star},\bm{R}_{\star}\right)\leq\varepsilon_{0}\sigma^{\star}_{r}<\sigma^{\star}_{r}$ provided $0\leq\alpha\leq\frac{\varepsilon_{0}}{50c_{s}\kappa\mu r\sqrt{r}}$. Thus, for $c\geq 1+\varepsilon_{0}$, by Lemma 7 we have
@@
and $\max\{\left\lVert{\bm{L}}_{0}{\bm{R}}_{0}^{*}\right\lVert_{2,\infty},\left\lVert{\bm{R}}_{0}{\bm{L}}_{0}^{*}\right\lVert_{2,\infty}\}\leq c\sqrt{\frac{c_{s}\mu r}{n}}\sigma^{\star}_{1}$. This completes the proof. ∎
5.4 Proof of Theorems 2 and 3 (local convergence)
In this subsection, based on the probabilistic events in Section 5.1, we prove several inequalities that are crucial for the proof of our main theorems. The proof follows a similar route established in [36], but the techniques and details are quite involved due to the extra Hankel structure, as well as the simultaneous presence of outliers and missing data.
@@ 5.4 Proof of Theorems 2 and 3 (local convergence)
$\max\{\left\lVert\bm{\Delta}_{{\bm{L}}_{k}}\bm{\Sigma}_{\star}^{-1/2}\right\lVert_{2},\left\lVert\bm{\Delta}_{{\bm{R}}_{k}}\bm{\Sigma}_{\star}^{-1/2}\right\lVert_{2}\}$ $\leq\varepsilon_{0},$ (28a)
$\max\left\{\left\lVert\bm{L}_{\natural}\bm{\Sigma}_{\star}^{1/2}\right\lVert_{2,\infty},\left\lVert\bm{R}_{\natural}\bm{\Sigma}_{\star}^{1/2}\right\lVert_{2,\infty}\right\}$ $\leq\frac{c}{1-\varepsilon_{0}}\sqrt{\frac{c_{s}\mu r}{n}}\sigma^{\star}_{1},$ (28b)
$\max\left\{\left\lVert\bm{L}_{\natural}\bm{\Sigma}_{\star}^{-1/2}\right\lVert_{2,\infty},\left\lVert\bm{R}_{\natural}\bm{\Sigma}_{\star}^{-1/2}\right\lVert_{2,\infty}\right\}$ $\leq\frac{c\kappa}{1-\varepsilon_{0}}\sqrt{\frac{c_{s}\mu r}{n}},$ (28c)
$\left\lVert\bm{R}_{\natural}\left(\bm{R}_{\natural}^{*}\bm{R}_{\natural}\right)^{-1}\bm{\Sigma}_{\star}^{1/2}\right\lVert_{2}$ $\leq\frac{1}{1-\varepsilon_{0}}.$ (28d)
Proof.
@@
then by $\left\lVert{\bm{L}}_{k}{\bm{R}}_{k}^{*}\right\lVert_{2,\infty}\geq\sigma_{r}(\bm{R}_{\natural}\bm{\Sigma}_{\star}^{-1/2})\left\lVert\bm{L}_{\natural}\bm{\Sigma}_{\star}^{1/2}\right\lVert_{2,\infty}$ we have $\left\lVert\bm{L}_{\natural}\bm{\Sigma}_{\star}^{1/2}\right\lVert_{2,\infty}\leq\frac{\left\lVert{\bm{L}}_{k}{\bm{R}}_{k}^{*}\right\lVert_{2,\infty}}{1-\varepsilon_{0}}$. Similarly, we have $\left\lVert\bm{R}_{\natural}\bm{\Sigma}_{\star}^{1/2}\right\lVert_{2,\infty}\leq\frac{\left\lVert{\bm{L}}_{k}{\bm{R}}_{k}^{*}\right\lVert_{2,\infty}}{1-\varepsilon_{0}}$ and conclude (28b). By using $\left\lVert{\bm{A}}{\bm{B}}\right\lVert_{2,\infty}\leq\left\lVert{\bm{A}}\right\lVert_{2,\infty}\left\lVert{\bm{B}}\right\lVert_{2}$, (28c) follows directly from (28b). (28d) follows from [36, Lemma 25] and (28a). ∎
Lemma 9.
For all matrices $({\bm{L}}_{k},{\bm{R}}_{k})\in\mathcal{E}(\varepsilon_{0}\sigma^{\star}_{r},c\sqrt{\frac{c_{s}\mu r}{n}}\sigma^{\star}_{1})$, it holds that
@@
$\leq\sqrt{p}(1+3\sqrt{\varepsilon_{0}})\left(\left\lVert\bm{\Delta}_{{\bm{L}}_{k}}\bm{\Sigma}_{\star}^{1/2}\right\lVert_{\mathrm{F}}+\left\lVert\bm{\Delta}_{{\bm{R}}_{k}}\bm{\Sigma}_{\star}^{1/2}\right\lVert_{\mathrm{F}}\right)$
@@
where the first step follows from $\left\lvert\left\langle\Pi_{\Omega}(\bm{u}+\bm{v}),\bm{u}+\bm{v}\right\rangle\right\rvert^{\frac{1}{2}}\leq\left\lvert\left\langle\Pi_{\Omega}\bm{u},\bm{u}\right\rangle\right\rvert^{\frac{1}{2}}+\left\lvert\left\langle\Pi_{\Omega}\bm{v},\bm{v}\right\rangle\right\rvert^{\frac{1}{2}}$ for $\bm{u},\bm{v}\in\mathbb{C}^{m}$, the second inequality follows from Lemma 5, and in the last inequality we have used $\left\lVert\bm{\Delta}_{{\bm{L}}_{k}}\bm{\Sigma}_{\star}^{1/2}{\bm{V}}_{\star}^{*}\right\lVert_{\mathrm{F}}=\left\lVert\bm{\Delta}_{{\bm{L}}_{k}}\bm{\Sigma}_{\star}^{1/2}\right\lVert_{\mathrm{F}}$. For the last term, provided $m\geq 6\varepsilon_{0}^{-2}C_{\varepsilon_{0},\kappa}^{2}c_{s}^{2}\mu^{2}r^{2}\log n$, we have
$~{}\frac{1}{p}\left\lvert\left\langle\Pi_{\Omega}\mathcal{G}^{*}\left(\bm{\Delta}_{{\bm{L}}_{k}}\bm{\Delta}_{{\bm{R}}_{k}}^{*}\right),\mathcal{G}^{*}\left(\bm{\Delta}_{{\bm{L}}_{k}}\bm{\Delta}_{{\bm{R}}_{k}}^{*}\right)\right\rangle\right\rvert$
@@
$~{}+\sqrt{\frac{24n\log n}{p}}\left\lVert\bm{\Delta}_{{\bm{L}}_{k}}\bm{\Sigma}_{\star}^{1/2}\right\lVert_{2,\infty}\left\lVert\bm{\Delta}_{{\bm{R}}_{k}}\bm{\Sigma}_{\star}^{-1/2}\right\lVert_{2,\infty}\left\lVert\bm{\Delta}_{{\bm{L}}_{k}}\bm{\Sigma}_{\star}^{1/2}\right\lVert_{\mathrm{F}}\left\lVert\bm{\Delta}_{{\bm{R}}_{k}}\bm{\Sigma}_{\star}^{-1/2}\right\lVert_{\mathrm{F}}$
@@
where the first inequality follows from Lemma 4, the second inequality follows from the inequalities $\left\lVert\bm{\Delta}_{{\bm{R}}_{k}}\bm{\Sigma}_{\star}^{-1/2}\right\lVert_{\mathrm{F}}=\left\lVert\bm{\Delta}_{{\bm{R}}_{k}}\bm{\Sigma}_{\star}^{1/2}\bm{\Sigma}_{\star}^{-1}\right\lVert_{\mathrm{F}}\leq\frac{1}{\sigma^{\star}_{r}}\left\lVert\bm{\Delta}_{{\bm{R}}_{k}}\bm{\Sigma}_{\star}^{1/2}\right\lVert_{\mathrm{F}}$, and (by (28c))
$\begin{split}\left\lVert\bm{\Delta}_{{\bm{L}}_{k}}\bm{\Sigma}_{\star}^{-\frac{1}{2}}\right\lVert_{2,\infty}&\leq\left\lVert\bm{L}_{\natural}\bm{\Sigma}_{\star}^{-\frac{1}{2}}\right\lVert_{2,\infty}+\left\lVert\bm{L}_{\star}\bm{\Sigma}_{\star}^{-\frac{1}{2}}\right\lVert_{2,\infty}\leq\left(1+\frac{c\kappa}{1-\varepsilon_{0}}\right)\sqrt{\frac{c_{s}\mu r}{n}},\cr\left\lVert\bm{\Delta}_{{\bm{L}}_{k}}\bm{\Sigma}_{\star}^{\frac{1}{2}}\right\lVert_{2,\infty}&\leq\left\lVert\bm{L}_{\natural}\bm{\Sigma}_{\star}^{\frac{1}{2}}\right\lVert_{2,\infty}+\left\lVert\bm{L}_{\star}\bm{\Sigma}_{\star}^{\frac{1}{2}}\right\lVert_{2,\infty}\leq\left(1+\frac{c}{1-\varepsilon_{0}}\right)\sqrt{\frac{c_{s}\mu r}{n}}\sigma^{\star}_{1}.\end{split}$ (29)
@@
Putting all the bound of terms together, we conclude the lemma. ∎
Lemma 10.